OpenStudy (anonymous):

use quadratic formula: x^2 - 15 = 3x PLEASE HELP!!!

OpenStudy (anonymous):

first set = 0 via \[x^2-3x-15=0\] \[(x-5)(x+3)=0\] \[x-5=0\] \[x=5\] or \[x+3=0\] \[x=-3\]

OpenStudy (anonymous):

oops sorry quadratic formula! hello daniel

OpenStudy (anonymous):

my answer is wrong wrong wrong

OpenStudy (anonymous):

\[Quadratic formula: x = (-b \pm \sqrt{b^2-4ac})/2a\] a = 1 b = -3 c = -15 \[x = (3\pm \sqrt{69})/2 \]

OpenStudy (anonymous):

i will bow out and let daniel reply. his answer will be correct. this one does not factor over integer.

OpenStudy (anonymous):

His answer "will be" correct? :D I love you satellite :D

OpenStudy (anonymous):

feeling is mutual (in a platonic kinda way)

OpenStudy (anonymous):

Thanks Guyss =)

OpenStudy (anonymous):

Who the heck is Platonic? Just kidding :D

OpenStudy (anonymous):

a solid fellow

OpenStudy (anonymous):

^ I was kidding :D

OpenStudy (anonymous):

What's The Final Answer though? i Don't Have A Calculator =/

OpenStudy (anonymous):

x = 5.65331193 or x = -2.65331193